OpenStudy (anonymous):

a + 3/4 = 7 1/5 what does a equal???

OpenStudy (anonymous):

An easy way is just to put it in fraction first and then finding a

OpenStudy (anonymous):

Not fraction damn.... I meant decimals.

OpenStudy (anonymous):

how would i convert 7 1/5 into a fraction???

OpenStudy (anonymous):

1/5 = 0.2 so 7.2

OpenStudy (anonymous):

oh okay thank you. :P

OpenStudy (anonymous):

No problem ;)

OpenStudy (anonymous):

7 and 1/5 - 3/4 = 6 and 9/20

OpenStudy (anonymous):

g + 4.9 = 8.3 how do I get the answer???

OpenStudy (anonymous):

Substract 4.9 from both sides to get g by itself.

OpenStudy (anonymous):

thanks.